# Semi-supervised Anomaly Detection via Adaptive Reinforcement Learning-Enabled Method with Causal Inference

### v1

@@ Abstract
Semi-supervised anomaly detection for guaranteeing the reliability of intelligent systems has received increasing attention. However, existing methods rely too much on data correlation and neglect causality, which can be misleading due to confounding factors and affect system reliability. Additionally, the current reinforcement learning anomaly detection methods can effectively identify known and unknown anomalies in environments with limited labeled samples. Despite its effectiveness, these methods still face several challenges, such as under-utilization of priori knowledge, lack of model flexibility, and insufficient reward feedback when interacting with the environment. To address the above problems, this paper innovatively constructs a counterfactual causal reinforcement learning model, termed Triple-Assisted Causal Reinforcement Learning Anomaly Detector (Tri-CRLAD). The model utilizes the causal inference mechanism to radically improve the performance of semi-supervised models and enhance the model’s ability to uncover anomaly data in the face of unknown or rare data. In addition, Tri-CRLAD features a triple decision support mechanism, namely, a sampling strategy based on historical similarity, an adaptive threshold smoothing adjustment strategy, and an adaptive decision reward mechanism. These mechanisms further enhance the flexibility and generalization ability of the model, enabling it to effectively respond to various complex and dynamically changing environments. Finally, Tri-CRLAD matches or exceeds the performance of 9 baseline methods across 7 diverse intelligent system datasets, including satellite systems, medical systems, and health systems. Moreover, anomaly detection stability was significantly improved by up to 23% with an extremely small number of known anomaly samples. Our code is available at https://github.com/Aoudsung/Tri-CRLAD/
@@ 1 Introduction
Anomaly detection, often referred to as the process of identifying deviations from the norm, plays a pivotal role across a multitude of intelligent systems, including but not limited to healthcare surveillance, satellite monitoring, cybersecurity, and credit fraud detection. [1, 2]. This field is highly sought after due to its inherent capability to enhance the reliability of the data collected, thereby contributing to the overall reliability and robustness of systems [3, 4, 5].
@@
However, since anomaly data is less than normal data, and the types of anomalies in real environments are complex and variable, which makes it difficult to label them completely.[6] Semi-supervised anomaly detection has received much attention from researchers because it requires less labeled data and can be combined with many unlabeled data streams to identify anomalies effectively. Hence, in intricate environments, the method’s design should be strong enough to develop an efficient detection model with a restricted number of anomaly data and a large number of normal samples. Moreover, it should be capable of generalizing and detecting unknown anomalies.
@@
Deep Reinforcement Learning (DRL) can effectively leverage its characteristics to meet the needs of semi-supervised anomaly detection methods [7, 8, 9]. The distinctive feature of DRL is its optimization of the decision-making process through balancing exploration and exploitation. Where exploration represents the ability of the agent to improve and refine its execution policy, and exploitation represents the ability of the agent to utilize the knowledge it has already learned to make decisions. Due to this characteristic, DRL has been widely used in several domains, such as autonomous driving and robot control [10, 11, 12]. For the semi-supervised anomaly detection problem, the DRL-based anomaly detection method learns efficiently with a small number of labeled anomaly samples and continuously explores and adapts to new anomaly patterns in real-time data streams.
@@
Figure 1: Visualization of the effects of causal inference and sampling strategies. Subfigures (a) and (b) display the original data features alongside those processed by causal inference, each visualized using the t-SNE method. Subfigures (c) and (d) depict the agent’s frequency of visits to the same point over identical step counts, employing distance-based and historical similarity-based sampling strategies. The x-axis denotes the number of visits, and the y-axis indicates the number of points receiving equal visitation. When the x value is lower, a higher y value indicates a better sampling strategy. Subfigures (a) and (b) demonstrate the efficacy of causally processed features in identifying anomaly samples. Meanwhile, subfigures (c) and (d) illustrate the proposed sampling mechanism’s ability to enable the agent to explore a broader range of sample points within the same step count, thus reducing repetitive sampling of identical points.
@@
Problem 1: Underutilization of prior knowledge. Much of the previous work has focused on learning by utilizing correlations in the data rather than by utilizing the intrinsic causal relationships between the data. However, as shown in Figure 1a, the raw data features may not represent the difference between abnormal and normal data well, implying that abnormal data are not well identified by correlations alone. This makes the model susceptible to the adverse effects of confounding information (e.g., noise dimensions) present in the data, which reduces the model’s reliability. On the other hand, the limitation imposed by the inadequate sampling method also restricts the utilization of prior knowledge. As shown in Figure 1c and 1d, when the x value is lower, a higher y value indicates a better sampling strategy. Figure 1c illustrates outcomes from the DAPLN method’s distance-based sampling strategy, which was applied to 100 samples [8]. This approach risks model overfitting by repeatedly selecting the same data points, potentially diminishing the model’s performance. In contrast, as shown in Figure 1d, our proposed sampling strategy then explores the entire dataset well.
@@
Problem 3: Current reward mechanisms in environments with anomalous interactions are restrictive and fail to furnish adequate feedback. These methods [8, 13] rely on rudimentary value systems for feedback, such as (-1, 0, 1). This mechanism is inflexible and does not allow for dynamic adjustment of rewards in response to the prevailing detection context. When the scenario involves a scarcity of labeled data instead of an abundance, the probability of the intelligence encountering an anomaly data point diminishes. Consequently, if the reward system is confined to a narrow spectrum of situations, it compels the agent to explore a greater variety of behaviors, extending the exploration phase and impairing the model’s learning efficiency.
This paper introduces the Tri-CRLAD model to address the above issues by developing a causal inference model comprising two stages, intra-point and inter-point modeling, accompanied by a triple decision support mechanism rooted in historical information. Intra-point modeling focuses on analyzing dimensional correlations within the data. In contrast, inter-point modeling utilizes counterfactual causal inference to extract the intrinsic causal feature, thereby augmenting the DRL agent’s effectiveness in applying the prior knowledge. In addition, the triple decision support mechanism skillfully integrates environmental changes into the decision-making process, enabling the agent to better adapt to changes in the ever-changing detection environments and thus effectively deal with the problem of semi-supervised anomaly detection for tabular data. In summary, this paper makes the following contributions:
@@
    Innovatively introduces causal reinforcement learning into semi-supervised anomaly detection, creating a model that leverages counterfactual causal inference to unearth the intrinsic causal feature in data, thereby improving the DRL agent’s utilization of prior knowledge and generalization capability.
@@
    Proposed a triple decision support mechanism based on historical information, encompassing a sampling strategy based on historical similarity, an adaptive threshold smoothing adjustment strategy, and an adaptive decision reward mechanism. These innovations collectively address the aforementioned challenges and substantially enhance the model’s training efficiency and generalization ability.
@@
    Demonstrating that the performance of Tri-CRLAD is on par with or surpasses existing state-of-the-art semi-supervised anomaly detection methods in various domain-specific anomaly detection datasets.
@@ 2.2 Deep Reinforcement learning
Deep reinforcement learning has demonstrated its power in many domains, such as game control, autonomous driving, and other more complex decision-making tasks [27, 28]. DRL is a learning approach in which a learning subject (called the "agent") learns how to complete a task or reach a goal by interacting with the environment. DRL’s essence lies in optimizing decision-making by balancing exploration, the agent’s capacity to refine the policy, and exploitation, the agent’s use of acquired knowledge. This involves navigating unknown environments to maximize long-term rewards. Notably, DRL has been adeptly applied to anomaly detection by some researchers [7, 8]. However, these algorithms still have a series of problems, such as the problem of the incentive mechanisms for incomplete feedback capacity(e.g., simply setting the decision reward to 0, 1, and -1), the problem of tendencies towards local optimality in [8], and limitations with fixed threshold settings in [7].
To address these issues, we proposed several measures. Illustrated in Figure 1d, the sampling strategy based on historical similarity aims to diversify the agent’s exposure to different data points, enhancing exploration efficiency and reducing repetitive sampling. Moreover, to heighten the accuracy of DRL in anomaly detection, we replace fixed threshold methods with an adaptive threshold smoothing strategy. Lastly, to refine the DRL’s decision-making accuracy, we propose a decision reward mechanism that makes it possible to adjust the value of the reward feedback according to the environment in which the agent is located. This strategy allows the agent to receive more comprehensive reward information rather than the incomplete reward information given in previous approaches. More in-depth analysis related to these strategies is presented in the ablation experiments section.
@@ 3 Problem Formulation
3.1 Anomaly Decision-making Interactive Environment
Our objective is to train an agent using a training dataset $D$ containing two initial training subdatasets: $D_{a}$, comprising labeled anomaly samples but smaller in size, and $D_{u}$, containing a larger volume of unlabeled samples. The agent’s task is to assign an anomaly score to each data point. The aim is to devise an anomaly scoring function, $\phi(\cdot):\mathbb{R^{n}}\rightarrow\mathbb{R}$, ensuring that for any anomaly data point $o_{i}$ and normal data point $o_{j}$, the condition $\phi(o_{i})>\phi(o_{j})$ holds true, where $o_{i},o_{j}\in D$.
@@ 3.1 Anomaly Decision-making Interactive Environment
  • Unknown Data Pool $U$: A data pool consisting of unlabeled data with detection.
@@ 3.2 Anomaly Detection Agent
The primary purpose of the ADIE is to select data for analysis from the sample pool using a predetermined sampling strategy and present them as states to the agent. Upon the agent’s decision-making and feedback to ADIE, ADIE must then provide appropriate reward information to the agent based on the reward strategy, facilitating updates to its scoring policy. To enhance the agent’s decision-making comprehensiveness, we employ the Soft Actor-Critic (SAC) reinforcement learning algorithm, suited for continuous action spaces [34]. SAC, an Actor-Critic-based algorithm, promotes exploration by incorporating entropy as an additional reward component, thereby encouraging the development of a robust anomaly scoring policy. In our study, the action space for the SAC agent ranges from $[0,1]$, with values closer to 1 indicating a higher likelihood of the tested sample being an anomaly. SAC’s objective is to maximize the combination of cumulative returns and entropy. Entropy, reflecting the policy’s randomness, spurs the algorithm to explore a broader range of behaviors. The objective function of SAC is represented by the equation:
@@
Where $\pi_{\phi}$ is the policy represented by a neural network that outputs a probability distribution of actions in a given state. $\rho_{\pi_{\phi}}$ is the distribution of the state and action under the policy $\pi_{\phi}$, $R(o_{t},a_{t})$ is the reward gained from taking action $a_{t}$ in the state $o_{t}$, $\gamma$ is the discount factor, $\mathcal{H}$ is the entropy, $\alpha$ is the entropy coefficient, which is used to control the importance of entropy in the importance in the objective function.
@@ 3.3 Counterfactual Causal Inference
Traditional reinforcement learning primarily relies on Correlation rather than Causation. However, relying only on correlation in complex and dynamic environments may lead to poor performance or incorrect decisions. In contrast, in Causal Reinforcement Learning, the model understands the causal relationships between variables through causal inference, which provides insight into the true impact of a particular action on the environment, which can help to make more effective and adaptive decisions. This approach sheds light on the genuine impact of actions on the environment, facilitating more effective and adaptive decision-making. Within causal inference, random variables in a causal graph $G$ (a directed acyclic graph) are typically categorized into two groups: a set of covariates $X$, represented as $\{X_{1},\ldots,X_{n}\}$, and an outcome variable $Y$. To estimate causal effects, an intervention operation, denoted as $\text{do}(X_{j}=x)$, is commonly utilized, setting the covariate $X_{j}$ to a specific value $x$. The causal effect of $\text{do}(X_{j}=x)$ is then estimated using the backdoor formulation.
@@ 4 The proposed Model
4.1 Overall structure of the Tri-CRLAD
Figure 3: The overall structure of the Tri-CRLAD. Tri-CRLAD comprises three principal components: the Causal Feature Extraction module, the DRL agent based on the SAC algorithm, and the ADIE module.
This section provides a comprehensive description of the semi-supervised anomaly detection model we developed, named Tri-CRLAD, to ensure the reliability of collected data. As depicted in Figure 3, the model’s overall structure and data flow are illustrated. Tri-CRLAD comprises three principal components: the Causal Feature Extraction module, the DRL agent based on the SAC algorithm, and the ADIE module. The process unfolds as follows: the Causal Feature Extraction module extracts the intrinsic causal features from data points sourced from ADIE, and these features (States) are then fed into the SAC. SAC, upon receiving these states, makes decisions (Actions) and interacts with ADIE, which in turn provides feedback (Rewards) based on the decisions to guide its scoring policy update. Next, we’ll delve into the various interaction mechanisms designed specifically for ADIE.

### v2

@@ Abstract
Semi-supervised anomaly detection for sensor signals is critical in ensuring system reliability in smart manufacturing. However, existing methods rely heavily on data correlation, neglecting causality and leading to potential misinterpretations due to confounding factors. Moreover, while current reinforcement learning-based methods can effectively identify known and unknown anomalies with limited labeled samples, these methods still face several challenges, such as under-utilization of priori knowledge, lack of model flexibility, and deficient reward feedback during environmental interactions. To address the above problems, this paper innovatively constructs a counterfactual causal reinforcement learning model, termed Triple-Assisted Causal Reinforcement Learning Anomaly Detector (Tri-CRLAD). The model leverages causal inference to extract the intrinsic causal feature in data, enhancing the agent’s utilization of prior knowledge and improving its generalization capability. In addition, Tri-CRLAD features a triple decision support mechanism, including a sampling strategy based on historical similarity, an adaptive threshold smoothing adjustment strategy, and an adaptive decision reward mechanism. These mechanisms further enhance the flexibility and generalization ability of the model, enabling it to effectively respond to various complex and dynamically changing environments. Experimental results across seven diverse sensor signal datasets demonstrate that Tri-CRLAD outperforms nine state-of-the-art baseline methods. Notably, Tri-CRLAD achieves up to a 23% improvement in anomaly detection stability with minimal known anomaly samples, highlighting its potential in semi-supervised anomaly detection scenarios. Our code is available at https://github.com/Aoudsung/Tri-CRLAD/.
@@ 1 Introduction
In the era of smart manufacturing, growing sensing and computing capabilities have facilitated the development and monitoring of sensor signals in engineering and the application of anomaly detection. Anomaly detection, often referred to as the process of identifying deviations from the norm, plays a pivotal role in a multitude of sensor-based monitoring processes, including but not limited to healthcare surveillance, satellite monitoring, cybersecurity, and wind turbine monitoring [1, 2, 3]. This field is highly sought after due to its inherent capability to enhance the reliability of the data collected, thereby contributing to the overall reliability and robustness of systems [4, 5, 6].
@@
However, since anomaly data is less than normal data, and the types of anomalies in real environments are complex and variable, which makes it is difficult to label them completely [7] in real scenarios. Semi-supervised anomaly detection has received much attention from researchers because it requires less labeled data and can be combined with many unlabeled data streams to identify anomalies effectively. Hence, in intricate industry environments, the method’s design should be strong enough to develop an efficient detection model with a restricted number of anomaly data and a large number of normal samples. Moreover, it should be capable of generalizing and detecting unknown anomalies.
@@
Deep Reinforcement Learning (DRL) can effectively leverage its characteristics to meet the needs of semi-supervised anomaly detection methods [8, 9, 10]. The distinctive feature of DRL is its optimization of the decision-making process through balancing exploration and exploitation. Where exploration represents the ability of the agent to improve and refine its execution policy, and exploitation represents the ability of the agent to utilize the knowledge it has already learned to make decisions. Due to this characteristic, DRL has been widely used in several domains, such as autonomous driving, Cyber-Physical Energy System Operations and Maintenance, and robot control [11, 12, 13, 14]. For the semi-supervised anomaly detection problem, the DRL-based anomaly detection method learns efficiently with a small number of labeled anomaly samples and continuously explores and adapts to new anomaly patterns in real-time data streams.
@@
Figure 1: Visualization of the effects of causal inference and sampling strategies. Subfigures 1a and 1b present the original versus causally inferred data features using t-SNE. Subfigures 1c and 1d depict the agent’s frequency of visits to the same point over identical step counts, employing distance-based and historical similarity-based sampling strategies. The x-axis represents visit frequency, and the y-axis shows the count of points with equal visits. A higher y value at a lower x value suggests a superior sampling strategy. Subfigures 1a and 1b highlight the efficacy of causally processed features in identifying anomaly samples. Meanwhile, subfigures 1c and 1d demonstrate the proposed sampling strategy’s effectiveness in facilitating a broader exploration of data points within a fixed step count, reducing the redundancy of revisiting identical points.
@@
Problem 1: Underutilization of prior knowledge. Much of the previous work has focused on learning by utilizing correlations in the data rather than by utilizing the intrinsic causal relationships between sensor signal data. However, as shown in Figure 1a, the raw data features may not represent the difference between abnormal and normal data well, implying that abnormal data are not well identified by correlations alone. This makes the model susceptible to the adverse effects of confounding information (e.g., noise dimensions) present in the data, which reduces the model’s reliability. On the other hand, the limitation imposed by the inadequate sampling method also restricts the utilization of prior knowledge. As shown in Figure 1c and 1d, when the x value is lower, a higher y value indicates a better sampling strategy. Figure 1c illustrates outcomes from the DAPLN method’s distance-based sampling strategy, which was applied to 100 samples [9]. This approach risks model overfitting by repeatedly selecting the same data points, potentially diminishing the model’s performance. In contrast, as shown in Figure 1d, our proposed sampling strategy then explores the entire dataset well.
@@
Problem 3: Current reward mechanisms in environments with anomalous interactions are restrictive and fail to furnish adequate feedback. These methods [9, 15] rely on a rudimentary value feedback mechanism, such as (-1, 0, 1). This mechanism is inflexible and does not allow for dynamic adjustment of rewards in response to the prevailing sensor signal context. When the scenario involves a scarcity of labeled data instead of an abundance, the probability of the intelligence encountering an anomaly data point diminishes. Consequently, if the reward system is confined to a narrow spectrum of situations, it compels the agent to explore a greater variety of behaviors, extending the exploration phase and impairing the model’s learning efficiency.
In order to address the above problems, this paper introduces the Tri-CRLAD model. Tri-CRLAD developed a causal inference model comprising two stages, intra-point and inter-point modeling, accompanied by a triple decision support mechanism rooted in historical information. Intra-point modeling focuses on analyzing dimensional correlations within the sensor signal data. In contrast, inter-point modeling utilizes counterfactual causal inference to extract the intrinsic causal feature, thereby augmenting the DRL agent’s effectiveness in applying the prior knowledge. In addition, the triple decision support mechanism skillfully integrates environmental changes into the decision-making process, enabling the agent to better adapt to changes in the ever-changing detection environments and thus effectively deal with the problem of semi-supervised anomaly detection for tabular sensor signal data. In summary, this paper makes the following contributions:
@@
    Innovatively integrates causal reinforcement learning into semi-supervised anomaly detection, developing the Tri-CRLAD that employs counterfactual causal inference to identify intrinsic causal features in the data. This procession enhances the DRL agent’s use of prior knowledge and its generalization capability.
@@
    Introduces a comprehensive triple decision support mechanism based on historical information. This mechanism includes a sampling strategy grounded in historical similarity, an adaptive threshold smoothing adjustment strategy, and an adaptive decision reward mechanism. These innovations collectively tackle the identified challenges and significantly improve the Tri-CRLAD’s training efficiency and generalization ability.
@@
    Extensive experiments demonstrate that the Tri-CRLAD achieves performance that meets or exceeds the current state-of-the-art semi-supervised anomaly detection methods across multiple sensor signal datasets.
@@ 2.2 Deep Reinforcement learning
Deep reinforcement learning has demonstrated its power in many domains, such as game control, autonomous driving, and other more complex decision-making tasks [29, 30]. DRL is a learning approach in which a learning subject (called the "agent") learns how to complete a task or reach a goal by interacting with the environment. DRL’s essence lies in optimizing decision-making by balancing exploration, the agent’s capacity to refine the policy, and exploitation, the agent’s use of acquired knowledge. This involves navigating unknown environments to maximize long-term rewards. Notably, DRL has been adeptly applied to anomaly detection by some researchers [8, pang2021toward,ijcai2023p19]. However, these algorithms still have a series of problems, such as the problem of the incentive mechanisms for incomplete feedback capacity(e.g., simply setting the decision reward to 0, 1, and -1), the problem of tendencies towards local optimality in [9], and limitations with fixed threshold settings in [8].
To address these issues, we proposed several measures. Illustrated in Figure 1d, the sampling strategy based on historical similarity aims to diversify the agent’s exposure to different data points, enhancing exploration efficiency and reducing repetitive sampling. Moreover, to heighten the accuracy of DRL in anomaly detection, we replace fixed threshold methods with an adaptive threshold smoothing strategy. Lastly, to refine the DRL’s decision-making accuracy, we propose a decision reward mechanism that makes it possible to adjust the value of the reward feedback according to the environment in which the agent is located. This strategy allows the agent to receive more comprehensive reward information rather than the incomplete reward information given in previous approaches. More in-depth analysis related to these strategies is presented in the ablation study section.
@@ 3 Problem Formulation
3.1 Anomaly Decision-making Interactive Environment for Sensor Signals
Ensuring the reliability of sensor signals in smart manufacturing is critical. Anomalies in sensor data can indicate potential faults or inefficiencies. Our objective is to train an agent using a training dataset $D$, which contains two initial training subdatasets: $D_{a}$, comprising labeled anomaly sensor signal data but smaller in size, and $D_{u}$, containing a larger volume of unlabeled sensor signal data. The agent’s task is to assign an anomaly score to each sensor signal data point. The aim is to devise an anomaly scoring function, $\phi(\cdot):\mathbb{R^{n}}\rightarrow\mathbb{R}$, ensuring that for any anomaly data point $o_{i}$ and normal data point $o_{j}$, the condition $\phi(o_{i})>\phi(o_{j})$ holds true, where $o_{i},o_{j}\in D$.
@@ 3.1 Anomaly Decision-making Interactive Environment for Sensor Signals
  • Unknown Data Pool $U$: A data pool consisting of unlabeled data pending detection.
@@ 3.2 Anomaly Detection Agent
The primary purpose of the ADIE is to select sensor signal data for analysis from the sample pool using a predetermined sampling strategy and present them as states to the agent. Upon the agent’s decision-making and feedback to ADIE, ADIE must then provide appropriate reward information to the agent based on the reward strategy, facilitating updates to its scoring policy. To enhance the agent’s decision-making comprehensiveness, we employ the Soft Actor-Critic (SAC) reinforcement learning algorithm, suited for continuous action spaces. SAC, an Actor-Critic-based algorithm, promotes exploration by incorporating entropy as an additional reward component, thereby encouraging the development of a robust anomaly scoring policy. In our study, the action space for the SAC agent ranges from $[0,1]$, with values closer to 1 indicating a higher likelihood of the tested sample being an anomaly. SAC’s objective is to maximize the combination of cumulative returns and entropy. Entropy, reflecting the policy’s randomness, spurs the algorithm to explore a broader range of behaviors. The objective function of SAC is represented by the equation:
@@
Where $\pi_{\phi}$ is the policy represented by a neural network that outputs a probability distribution of actions in a given state. $\rho_{\pi_{\phi}}$ is the distribution of the state and action under the policy $\pi_{\phi}$, $R(o_{t},a_{t})$ is the reward gained from taking action $a_{t}$ in the state $o_{t}$, $\gamma$ is the discount factor, $\mathcal{H}$ is the entropy, $\alpha$ is the entropy coefficient, which is used to control the importance of entropy in the objective function.
@@ 3.3 Counterfactual Causal Inference
Traditional reinforcement learning primarily relies on Correlation rather than Causation. However, relying only on correlation in complex and dynamic environments may lead to poor performance or incorrect decisions. In contrast, in Causal Reinforcement Learning, the model understands the causal relationships between variables through causal inference, which provides insight into the true impact of a particular action on the environment. This approach sheds light on the genuine impact of actions on the environment, facilitating more effective and adaptive decision-making. Within causal inference, random variables in a causal graph $G$ (a directed acyclic graph) are typically categorized into two groups: a set of covariates $X$, represented as $\{X_{1},\ldots,X_{n}\}$, and an outcome variable $Y$. To estimate causal effects, an intervention operation, denoted as $\text{do}(X_{j}=x)$, is commonly utilized, setting the covariate $X_{j}$ to a specific value $x$. The causal effect of $\text{do}(X_{j}=x)$ is then estimated using the backdoor formulation.
@@ 4 The proposed Model
4.1 Overall structure of the Tri-CRLAD for Sensor Signals
Figure 3: The overall structure of the Tri-CRLAD. To ensure the reliability of the collected sensor signal data, Tri-CRLAD comprises three principal components: the Causal Feature Extraction module, the DRL agent based on the SAC algorithm, and the ADIE module.
To further illustrate how we ensure the reliability of the collected sensor signal data, this section provides a comprehensive description of the semi-supervised anomaly detection model we developed, named Tri-CRLAD. As depicted in Figure 3, the model’s overall structure and data flow are illustrated. Tri-CRLAD comprises three principal components: the Causal Feature Extraction module, the DRL agent based on the SAC algorithm, and the ADIE module. The process unfolds as follows: the Causal Feature Extraction module extracts the intrinsic causal features from data points sourced from ADIE, and these features (States) are then fed into the SAC. SAC, upon receiving these states, makes decisions (Actions) and interacts with ADIE, which in turn provides feedback (Rewards) based on the decisions to guide its scoring policy update. Next, we’ll delve into the various interaction mechanisms designed specifically for ADIE.